OpenStudy (gabylovesyou):

The number of Math homework questions is seven more than one−half the number of History homework questions. There are 15 Math homework questions. Which of the following equations would you use to find the number of History homework questions? Note: h represents the number of questions on your History homework.

OpenStudy (gabylovesyou):

15/2+7=h 15+7/2=h h+7/2=15 1/2h+7=15

OpenStudy (anonymous):

are you trying to find h yes??

OpenStudy (gabylovesyou):

yup

OpenStudy (anonymous):

 do you follow so far? :D